# algorithmic

### 10 Dec 2023 v3

@@ 4.2 Lie Algebra Convolutions
SE(2) 10 Almost Equivariant G-CNN 92.45 $\mathbf{0.0363\pm 0.0004}$ $\mathbf{0.5571\pm 2.1730}$
E(2) 10 E2CNN 92.81 $0.0349\pm 0.0001$ $3.5987\pm 2.8203$
Residual Pathway Prior 85.20 $0.0350\pm 0.0001$ $14.4018\pm 26.8171$
N/A Approximately Equivariant G-CNN 85.51 $0.1349\pm 0.1414$ $1.4893\pm 1.8695$
T(2) N/A Standard CNN 92.05 $0.0354\pm 0.0009$ $0.6573\pm 1.0565$
Table 1: Rot-MNIST classification accuracies and RMSE prediction errors for pendulum trajectory prediction. The first column gives the Lie group with respect to which equivariance or almost equivariance (depending on the model) is imposed. The second column gives the number of Lie group/algebra elements sampled by the model, if applicable.
@@ 5.3 Smoke Plume
Finally, we test our model on an almost equivariant prediction task. The dataset we use is the smoke plume dataset of Wang et al. (2022a) consisting of $64\times 64$ 2D velocity vector fields of smoke simulations with different initial conditions and external forces, all generated using the PDE simulation framework, PhiFlow (Holl et al., 2020). Specifically, we use the subset of the data that features rotational almost equivariance. As stated in Wang et al. (2022a), “both the inflow location and the direction of the buoyant forces possess a perfect rotation symmetry with respect to the $C_{4}$ group, but the buoyancy factor varies with the inflow positions to break the rotational symmetry.” All models are trained to predict the raw velocity fields at the next time step given the raw velocity fields at the previous timestep as input.
@@
Group Num Samples Model Future (RMSE) Domain (RMSE)
SE(2) 4 Almost Equivariant G-CNN 1.18 0.78
E(2) 4 E2CNN 1.05 0.76
Residual Pathway Prior 0.96 0.83
Steerable Approximately Equivariant G-CNN 0.80 0.67
T(2) N/A Standard CNN 1.21 1.10
Table 2: Prediction RMSE on simulated smoke plume velocity fields with almost rotational symmetry. The results for the baseline methods are taken from Wang et al. (2022a) and compared against our Almost Equivariant GCNN. As stated in Wang et al. (2022a), Future prediction involves testing on data that lies in the future of the training data. Domain prediction involves training and test data that are from different spatial domains.

### 02 Feb 2024 v4

@@ 2.2 Almost Equivariance
2.3 Notions of Approximate Equivariance
It’s worthwhile to note that there are multiple notions of approximate, partial, and soft equivariance, only some of which we explicitly address in this work.
The first type occurs when we only have partially observed data, for example, a single pose of a 3D object captured in a 2D image or an object occlusion in computer vision. Wang et al. (2023) refer to this as extrinsic equivariance in that applying a group transformation to an in-distribution data point transforms it to an out-of-distribution data point. This type of partial equivariance is often addressed via data augmentation. We do not explicitly test our approach in this setting.
The second type occurs when we have noise in data that breaks equivariance. This is one setting we explicitly address.
The third type occurs when we have data that naturally exhibits almost equivariance. For example, data sampled from vector fields and PDEs governing natural physical processes often exhibit this quality. This is another setting we exxplicitly address.
Finally, there is what Wang et al. (2023) call incorrect equivariance. This occurs when applying a group transformation to a data point qualitatively and quantitatively changes its label. For example, rotating the digit 6 by 180 degrees turns it into the digit 9 and vice versa. We do not explicitly address this in our method, but our model performs competitively on the Rot-MNIST classification task, indicating that it has the capability of accounting for incorrect equivariances in its modeling.
@@ 4.2 Lie Algebra Convolutions
SE(2) 10 Almost Equivariant G-CNN $\mathbf{92.05\pm 0.27}$ $\mathbf{0.0363\pm 0.0004}$ $\mathbf{0.5571\pm 2.1730}$
E(2) 10 E2CNN $91.91\pm 0.22$ $0.0349\pm 0.0001$ $3.5987\pm 2.8203$
Residual Pathway Prior $85.20\pm 0.66$ $0.0350\pm 0.0001$ $14.4018\pm 26.8171$
N/A Approximately Equivariant G-CNN $84.99\pm 0.37$ $0.1349\pm 0.1414$ $1.4893\pm 1.8695$
T(2) N/A Standard CNN $85.95\pm 0.49$ $0.0354\pm 0.0009$ $0.6573\pm 1.0565$
Table 1: Rot-MNIST classification accuracies and RMSE prediction errors for pendulum trajectory prediction. The first column gives the Lie group with respect to which equivariance or almost equivariance (depending on the model) is imposed. The second column gives the number of Lie group/algebra elements sampled by the model, if applicable. Best results are bold-faced and second-best are colored gray.
@@ 5.3 Smoke Plume
Next, we test our model on an almost equivariant prediction task. The dataset we use is the smoke plume dataset of Wang et al. (2022a) consisting of $64\times 64$ 2D velocity vector fields of smoke simulations with different initial conditions and external forces, all generated using the PDE simulation framework, PhiFlow (Holl et al., 2020). Specifically, we use the subset of the data that features rotational almost equivariance. As stated in Wang et al. (2022a), “both the inflow location and the direction of the buoyant forces possess a perfect rotation symmetry with respect to the $C_{4}$ group, but the buoyancy factor varies with the inflow positions to break the rotational symmetry.” All models are trained to predict the raw velocity fields at the next time step given the raw velocity fields at the previous timestep as input.
@@
5.4 Jet Flow
Finally, we test on one more almost equivariant dataset. As described in Wang et al. (2022a), this dataset contains samples of 2D turbulent velocity fields taken from NASA multi-stream jets that were measured using time-resolved particle image velocimetry as described in Bridges & Wernet (2017). We follow the procedure described in Wang et al. (2022a), and “train and test on twenty-four $62\times 23$ sub-regions of jet flows.” Table 2 shows our results.
Group Num Samples Model Jet Flow (RMSE) Smoke Plume (RMSE)
Future Domain Future Domain
SE(2) 10 Almost Equivariant G-CNN $0.1931\pm 0.0012$ $0.2078\pm 0.0008$ 1.18 0.78
E(2) 10 E2CNN $0.1919\pm 0.0016$ $0.2131\pm 0.0023$ 1.05 0.76
10 Residual Pathway Prior $0.1947\pm 0.0066$ $0.2143\pm 0.0057$ 0.96 0.83
4 Steerable Approximately Equivariant G-CNN $\mathbf{0.1597\pm 0.0016}$ $\mathbf{0.1785\pm 0.0023}$ 0.80 0.67
T(2) N/A Standard CNN $0.2109\pm 0.0068$ $0.2218\pm 0.0008$ 1.21 1.10
Table 2: Prediction RMSE on simulated smoke plume velocity fields and jet flow 2D turbulent velocity fields with almost rotational symmetry. The results for the baseline methods are taken from Wang et al. (2022a) and compared against our Almost Equivariant G-CNN. As stated in Wang et al. (2022a), Future prediction involves testing on data that lies in the future of the training data. Domain prediction involves training and test data that are from different spatial domains. Best results are bold-faced and second-best are colored gray.
@@ A.2.1 Representation Theory
Representation theory seeks to extend the theory of linear algebra to groups (and more general objects, such as algebras) by associating to each group a representation, which is a homomorphism from the group to the space of linear operators on that group. As a simplification, we often just think of the representation as an association of an $n\times n$ matrix to each group element, in which case we have a matrix group.
@@ A.2.2 Differential Topology & Geometry, Lie Groups, and Lie Algebras
Smooth manifold theory extends the techniques of calculus to high-dimensional, non-Euclidean spaces and those without a preferred coordinate system. In layman’s terms, a smooth manifold is a mathematical object that is locally homeomorphic to $\mathbb{R}^{n}$ about every point and which has a smooth structure that allows one to perform operations from calculus such as differentiation and integration. More concretely, we equip a Hausdorff, second countable, and locally Euclidean topologyical space with a set of charts, $\{(U_{k},\varphi_{k})\}_{k=1}^{n}$ which consist of a neighborhood $U_{k}$ about each point and a homeomorphism $\varphi_{k}:U_{k}\to\mathbb{R}^{n}$. We then define transition maps, $\psi\circ\varphi^{-1}:\varphi(U\cap V)\to\psi(U\cap V)$ that allow us to move between charts. For smooth manifolds, we require that these charts are smoothly compatible, i.e. that either $U\cap V=\emptyset$ or $\psi\circ\varphi^{-1}$ is a diffeomorphism.